The Denver Post reported that, on average, a large shopping center had an incident of shoplifting caught by security 1.6 times every five hours. The shopping center is open from 10 A.M. to 9 P.M. (11 hours). Let r be the number of shoplifting incidents caught by security in an 11-hour period during which the center is open.

(a) Explain why the Poisson probability distribution would be a good choice for the random variable r.

Frequency of shoplifting is a common occurrence. It is reasonable to assume the events are independent.Frequency of shoplifting is a rare occurrence. It is reasonable to assume the events are independent.    Frequency of shoplifting is a common occurrence. It is reasonable to assume the events are dependent.Frequency of shoplifting is a rare occurrence. It is reasonable to assume the events are dependent.



What is λ? (Use 2 decimal places.)


(b) What is the probability that from 10 A.M. to 9 P.M. there will be at least one shoplifting incident caught by security? (Use 4 decimal places.)


(c) What is the probability that from 10 A.M. to 9 P.M. there will be at least three shoplifting incidents caught by security? (Use 4 decimal places.)


(d) What is the probability that from 10 A.M. to 9 P.M. there will be no shoplifting incidents caught by security? (Use 4 decimal places.)
Answer #1

a)

frequency of shoplifting is a rare occurrence. It is reasonable to assume the events are independent

λ =11*1.6/5=3.52 times every eleven hours

b)  probability that from 10 A.M. to 9 P.M. there will be at least one shoplifting incident caught by security

=P(X>=1) =1-P(X=0) =1-e-3.52*3.520/0! =0.9704

c)  probability that from 10 A.M. to 9 P.M. there will be at least three shoplifting incidents caught by security

=P(X>=3) =1-P(X<=2) =1-(P(X=0)+P(X=1)+P(X=2)) =1-(e-3.52*3.520/0!+e-3.52*3.521/1!+e-3.52*3.522/2!)

=1-(0.0296+0.1042+0.1834)=0.6828

d)probability that from 10 A.M. to 9 P.M. there will be no shoplifting incidents caught by security =P(X=0)

=e-3.52=0.0296